OpenStudy (anonymous):

4 Fractions Questions: 1) 4 2/7 x 2 1/5 2) -3 2/3 x (-2 1/4) 3) -4 2/5 / (-5) 4) 4 1/3 / 2 3/5

OpenStudy (anonymous):

The / in 3. and 4, is a divsion sign!!

OpenStudy (anonymous):

(4 + 2/7) (2 + 1/5) = 66/7 (-3 - 2/3) (-2 - 1/4) = 33/4 You should be able to do the last two.

OpenStudy (anonymous):

Thanks.